# Fix-Con: Automatic Fault Localization and Repair of Deep Learning Model Conversions between Frameworks

### v3

@@ V-A Fault Analysis Setup
V-B Fault Repair Setup
Once we extracted the order of layers from the fault analysis, we utilized the fault repair component of Fix-Con, applying the strategies in a greedy approach, as described in Section IV-B. We set a time limit of 2 hours for each conversion case. We also set a limit of three unsuccessful repair iterations, when repairs did not reduce output label discrepancy, as a stopping condition. This is to help make the experiments tractable. These settings can be customized as needed in our tool.
VI Results
Overall, Fix-Con was able to localize $755755$ potential faults in the structure and properties between Source and Target variants across all $1515$ erroneous conversion cases. An overview of the potential faults distribution can be found in Table II. From these $755755$ cases Fix-Con identified as bugs, it repaired $462462$ of them — completely repairing $1212$ out of $1515$ problematic model conversions, while significantly improving the performance of $22$ more, with the most distinctive case being improved from 48.948.948.948.9% output label discrepancy to 0.330.330.330.33%. Our system was unable to reduce the percentage further, as none of our strategies affected the model beyond this point.
VI-A Fault Localization
The faults localized were distributed across all six types within input-based and layer-based fault categories. Of the $755755$ faults localized, $462462$ of them (61.161.161.161.1% of total cases) resulted in reducing output label discrepancy between Source and Targetwhen repaired. The remaining localized faults did not result in any improvement after repair and are treated as false positives detected by the fault localization component. We demonstrate the distribution of faults detected and fixed across the different types in Table II and discuss each type in the context of fault localization next.
TABLE II: Number of faults localized versus those repaired (#Localisations/#Repairs) per model conversion case.
@@ 15 TF-to-TFLite 0/0 0/0 0/0 94/94 0/0 0/0
Input Dimensions (ID): Fix-Con identified four model cases where the input dimensions were different between Source and Target. We identified that input dimension errors had two primary causes: (1) erroneous configuration from the user, which neither the model converters nor the model itself raised as warnings or errors; (2) implicit changes in model input due to the converter implementation.
@@
We observed two cases, $99$ and $1313$ in Table II, where ID fault is a result of the first cause. Regarding the second cause, we detected two cases where the converter changed the model input. In addition to ID faults, we also detected differences in tensor shape and structure that we discuss as part of the next category.
@@
Tensor Shape and Structure (TSS): Fix-Con detected three cases with potential faults in the TSS category. The model converter introduced a Transpose node in these two cases to address the difference in ID so that subsequent nodes use inputs with the correct dimension. However, in these two cases getting to the correct input dimension with the Transpose node was not adequate. This occurred due to similar sized dimensions within tensors, like the third and fourth dimension within an example tensor with shape (1, 3, 224, 224), where it is possible to incorrectly transpose dimensions of the same size that result in a different structure from what is intended.
@@
In addition, in some model architectures (e.g., InceptionV3) the Transpose node is placed after a series of nodes that normalize data based on a specific input dimension. For instance, in InceptionV3, the Transpose node is placed after a series of Gather, Unsqueeze, Mul, and Add nodes. Fix-Con detects these differences across model graph and reports them during fault localization.
@@
Weights and Biases (WB): We observed $438438$ cases of differences in tensor values between weights and biases across Source and Target, with ResNet101 converted from Keras to TFLite containing $177177$ of them.
@@
Layer Hyperparameters (LH): Fix-Con detected $254254$ potential faults across $1212$ model conversion cases, with most of them concerning padding hyperparameters. Fix-Con also identified some cases where a potential hyperparameter fault was associated with a potential computation graph (CG) fault - for instance, specific convolutional layers presented potential faults related with padding hyperparameters and Pad nodes that directly preceded these convolutional layers.
@@
Computation Graph (CG): Our system identified $4747$ cases across $44$ conversion cases in which parts of the model computation graph presented discrepancies across Source and Target variants. All the cases observed were associated primarily with three operations: (1) setting a padding around a tensor, by utilizing Pad nodes; (2) performing batch normalization to a tensor by applying an element-by-element constant addition, and scaling using a BatchNorm node; and (3) redundant output nodes erroneously placed by the converter. From those $4747$ cases, Fix-Con successfully identified $44$ of those cases as actual bugs and repaired them, playing a crucial role for the repair of cases #$1111$ and $1313$.
VI-B Fault Repair
@@
PP: As shown in Table III, Fix-Con was able to repair cases #$11$, $33$, $66$, $77$, $99$ and $1010$, by determining and setting the correct preprocessing configuration. Given that the preprocessing is not a standard procedure across models, we identified that for these cases using the Source preprocessing instead of the one officially provided for the Target library fixed the result completely. Cases #$22$, $44$, $99$, and $1414$ were also affected partially by preprocessing selection, however additional steps were needed in order to repair them. The additional steps are explained in more detail in the categories below.
@@
ID: Among the $44$ faults localized as ID, Fix-Con repaired one case completely by solely adjusting the input dimensions of Target to match Source. For the remaining three cases, additional actions had to be performed, involving repair tensor shape and structure, as well as computation graph repair, as described in the next categories.
@@
TSS: Fix-Con was able to repair three cases related to problematic tensor shape and structure. In particular, cases $44$, $1313$, and $1414$ were related to problematic layers right after input. All three cases had different input dimensions between Source and Target, and the conversion tool had inserted additional Transpose nodes in order to adjust the input tensor shape so that it would be aligned with subsequent nodes. However, the transformation was done improperly, leading to severe performance degradation of up to 68.2168.2168.2168.21%. We identified all cases to be related to the onnxmltools [48] converter (part of the native ONNX codebase), when converting Keras to ONNX. Fix-Con automatically identified the presence of a redundant Transpose node following the input node, and repaired it by (1) adjusting the Target model input dimensions so that they correspond to those of Source, (2) altering its tensor transposition order (permutation) property to align with the new input dimensions, and (3) updating the properties of nodes connected to Transpose to ensure proper tensor handling, reflecting the change in Transpose. This modification in particular was vital to case #$1414$ where Fix-Con detected additional nodes (Gather, Unsqueeze, Add and Mul) preceding Transpose that apply input normalization. As part of the repair process, Fix-Con updated the axes properties of Gather and Unsqueeze nodes to align with the updated input dimension and the Transpose permutation in Target.
@@
WB: Cases #$22$, $55$, $88$, and $1515$ were related to parameter precision discrepancies between Source and Target. Although the discrepancies were negligible, the model behavior deviated across the cases from 3.123.123.123.12-9.79.79.79.7%. Fix-Con performed multiple iterations and managed to completely fix cases #$55$ and $1515$ while also significantly improving cases $22$ and $88$ by replacing the weights of Target with those of Sourcewhere differences were detected. We identified all cases to be related to the behavior and the setup of TFLiteConverter. The discrepancy after each repair step is depicted in Figure 4. At each repair step, Fix-Con adjusts the weights & biases of a suspicious layer in Target with those from Source, in the order determined by the Layer Activation Analysis (Section IV-A4), and evaluates the effects of the change by comparing the predictions of Source and Target.
@@
LH: From the $254254$ potential faults Fix-Con localized, only $1212$ of them were actively affecting the model. These hyperparameters were associated with the pads hyperparameter of Conv and MaxPool nodes (cases #$1111$ & $1313$), as well as with the axes of Gather and Unsqueeze nodes responsible for input processing (case #$1414$). Repairing Target entailed adjusting their values with those from Source.
@@
CG: Fix-Con identified $4747$ cases of graph-related potential faults. Of these, $44$ cases resulted in fault repairs that improved model performance (reduced output label discrepancy). The repairs played a crucial role in conversion cases #$1111$ and #$1313$ that initially presented discrepancies of 100100100100% and 35353535%, respectively. Our tool identified that, for case #$1111$, the converter had erroneously introduced multiple output nodes. In addition, three problematic padding nodes were found to affect the behavior of the models. Fix-Con identified such faults and completely repaired the models by (1) removing any additional output nodes added to the model by the converter, and (2) repairing problematic padding nodes.
@@
For all the false positive cases observed in LH and CG, we performed manual inspection and observed that the converters performed semantically equivalent computation graph conversions. For instance, in the erroneous conversion presenting the largest number of CG fault localizations – ResNet101, Keras to TFLite (case #$88$) – Source contained Pad nodes which were absent in Target. However, Target contained a pads hyperparameter in the next convolutional layer that had the same effect. Also, BatchNorm nodes of Source were replaced by the converter with Mul and Add nodes in Target  effectively applying the normalization in a two-step semantically equivalent manner. Fix-Con attempted to repair these scenarios by replacing them with the corresponding Source subgraph but these fixes had no effect on the number of output label discrepancies encountered.
@@
Figure 4: Model conversion cases that required an iterative weights and biases repair strategy with the percentage output label dissimilarity shown after each repair cycle (following preprocessing repair).
VI-C Fault Localization/Repair Effectiveness
@@
We consider effectiveness as the number of localized faults that could be repaired to improve Target prediction accuracy towards matching this of Source. We go through each fault category to determine the effectiveness of the fault localization component, as seen in the Total row in Table II. We find that Fix-Con had an 100100100100% efficacy rate (all potential fault occurrences detected could be repaired) for preprocessing (PP: 9/9999/99 / 9), input dimensions (ID: 4/4444/44 / 4) and tensor structure (TSS: 2/2222/22 / 2), as well as 98.198.198.198.1% efficacy for weights and biases (WB: 430/438430438430/438430 / 438). The weights and biases category had the most faults detected and repaired. Fix-Con had much less success on cases related to the computation graph (CG), with $44$ of the $4747$ potential faults detected being repaired (8.58.58.58.5%) while other cases when repaired did not result in reduction of output label discrepancy. The CG fault repairs, although few in number, were instrumental in completely fixing conversion cases #$1111$ and $1313$. The effectiveness for LH category was low (4444%), with only $1212$ out of the $254254$ identified potential faults worth repairing. Following manual inspection, we infer that this happens because: (1) graph transformations across Source and Target may result in syntactic difference but be semantically equivalent. For instance, a BatchNorm node is transformed in Mul and Add nodes. Although this is localized as a potential fault by Fix-Con  the repair employed by replacing Mul and Add nodes with BatchNorm will not reduce the output label discrepancy and is therefore rejected; (2) the LH difference detected is already mitigated by the converter by introducing additional nodes in the TargetḞor example, a pads (padding) hyperparameter in a Conv node on Source is replaced by a dominating Pad to the correspondent Conv node in target, leading to the same effect.
VI-D Repair Validity
@@
In order to verify our results, we used the ImageNetV2 [69] dataset consisting of 10K images which was built with the purpose of providing test data for ImageNet, as an independent validation source. We checked the discrepancies of the repaired Target models against Source, and verified that (1) all cases that were fully repaired using ILSVRC2017 also presented no discrepancies with ImageNetV2, and (2) partially repaired cases #$22$ and $88$ presented 0.280.280.280.28% and 0.660.660.660.66% discrepancies, lower than those observed in ILSVRC2017. The full comparison can be found in Table III. In addition, we used ImageNetV2 to verify that repaired models were giving consistent results against the dataset ground truths. Finally, we tested all $1010$ cases shown in Figure 1 presenting no discrepancies with ImageNetV2, to verify that no additional errors were detected.
VI-E Results Generalizability
@@
Our approach is applicable to any DL framework or conversion tool. In addition, it can be applied to any DL model architecture, including non-CNN models, only requiring the user to specify the core computational layers to be used in the comparison. We chose image recognition models in our experiments as they are more tractable and need less resources than other DL architectures like transformers, as our experiments involved $1515$ model conversion cases with numerous iterations and analysis within each. All the models selected are widely used, either as-is or as the base for extension of models deployed on more complex tasks (e.g., ResNet101 extended for object detection).
@@
In terms of the fault types supported by Fix-Con, they are based on a survey of issues encountered in discussion forums and empirical investigations (Table I). There may be other fault types not currently considered by Fix-Con. We are aware of this and we expect that Fix-Con will keep evolving to handle emerging issues and other fault categories in model conversions between DL frameworks.
@@
TABLE III: Experiments Overview: Discrepancies between Source and Target. before and after the repair process. for ILSVRC2017 and ImageNetV2 datasets.

### v4

@@ 5.1. Fault Analysis Setup
5.2. Fault Repair Setup
Once we extracted the order of layers from the fault analysis, we utilized the fault repair component of FetaFix, applying the strategies in a greedy approach, as described in Section 4.2. We set a time limit of 2 hours for each conversion case. We also set a limit of three unsuccessful repair iterations, when repairs did not reduce output label discrepancy, as a stopping condition. This is to help make the experiments tractable. These settings can be customized as needed in our tool.
6. Results
[…]
6.1. Fault Localization
The faults localized were distributed across all six types within input-based and layer-based fault categories. Of the $755755$ faults localized, $462462$ of them (61.161.161.161.1% of total cases) resulted in reducing output label discrepancy between Source and Target when repaired. For the remaining localized faults, FetaFix attempted to repair them by applying the respective repair strategy, but did not result in any improvement. Furthermore, they are considered as false positives detected by the fault localization component and are automatically discarded from the repair process. We demonstrate the distribution of faults detected and fixed across the different types in Table 2 and discuss each type in the context of fault localization next.
Overall, the faults we detected are new, as we detected no reported issues with relation to them in the respective converter repositories. We aim reporting the issues to the respective developers of the conversion tools.
Table 2. Number of faults localized versus repaired (#Localisations/#Repairs) per model conversion case.
@@ 15 TF-to-TFLite 0/0 0/0 0/0 94/94 0/0 0/0
Input Dimensions (ID): FetaFix identified four model cases where the input dimensions were different between Source and Target. We identified that input dimension errors had two primary causes: (1) erroneous configuration from the user, which neither the model converters nor the model itself raised as warnings or errors; (2) implicit changes in model input due to the converter implementation.
@@
We observed two cases, $99$ and $1313$ in Table 2, where ID fault is a result of the first cause. Regarding the second cause, we detected two cases where the converter changed the model input. In addition to ID faults, we also detected differences in tensor shape and structure that we discuss as part of the next category.
@@
Tensor Shape and Structure (TSS): FetaFix detected three cases with potential faults in the TSS category. The model converter introduced a Transpose node in these two cases to address the difference in ID, so that subsequent nodes use inputs with the correct dimension. However, in these two cases achieving the correct input dimension by using the Transpose node was not adequate. This occurred due to similarly sized dimensions within tensors, like the third and fourth dimension within an example tensor with shape (1, 3, 224, 224), where it is possible to incorrectly transpose dimensions of the same size that result in a different structure from what is intended.
@@
In addition, in some model architectures (e.g., InceptionV3) the Transpose node is placed after a series of nodes that normalize data based on a specific input dimension. For instance, in InceptionV3, the Transpose node is placed after a series of Gather, Unsqueeze, Mul, and Add nodes. FetaFix detects these differences across model graph and reports them during fault localization.
@@
Weights and Biases (WB): We observed $438438$ cases of tensor value differences in weights and biases between Source and Target, with ResNet101 converted from Keras to TFLite containing $177177$ of them.
@@
Layer Hyperparameters (LH): FetaFix detected $254254$ potential faults across $1212$ model conversion cases, with most of them concerning padding hyperparameters. FetaFix also identified some cases where a potential hyperparameter fault was associated with a potential computation graph (CG) fault - for instance, specific convolutional layers presented potential faults related with padding hyperparameters and Pad nodes that directly preceded these convolutional layers.
@@
Computation Graph (CG): FetaFix identified $4747$ cases across $44$ conversion cases in which parts of the model computation graph presented discrepancies across Source and Target variants. All the cases observed were associated primarily with three operations: (1) setting a padding around a tensor, by utilizing Pad nodes; (2) performing batch normalization to a tensor by applying an element-by-element constant addition, and scaling using a BatchNorm node; and (3) redundant output nodes erroneously placed by the converter. From those $4747$ cases, FetaFix successfully identified $44$ of them as actual bugs and repaired them, playing a crucial role for the repair of cases #$1111$ and $1313$.
6.2. Fault Repair
@@
PP: As shown in Table 3, FetaFix was able to repair cases #$11$, $33$, $66$, $77$, $99$ and $1010$, by determining and setting the correct preprocessing configuration. Given that the preprocessing is not a standard procedure across models, we identified that for these cases using the Source preprocessing instead of the one officially provided for the Target library fixed the result completely. Cases #$22$, $44$, $99$, and $1414$ were also affected partially by preprocessing selection, however additional steps were needed in order to repair them. The additional steps are explained in more detail in the categories below.
@@
ID: Among the $44$ faults localized as ID, FetaFix repaired one case completely by solely adjusting the input dimensions of Target to match Source. For the remaining three cases, additional actions had to be performed, involving repair tensor shape and structure, as well as computation graph repair, as described in the next categories.
@@
TSS: FetaFix was able to repair three cases related to problematic tensor shape and structure. In particular, cases $44$, $1313$, and $1414$ were related to problematic layers right after input. All three cases had different input dimensions between Source and Target, and the conversion tool had inserted additional Transpose nodes in order to adjust the input tensor shape so that it would be aligned with subsequent nodes. However, the transformation was done improperly, leading to severe performance degradation of up to 68.2168.2168.2168.21%. We identified all cases to be related to the onnxmltools (onn, 2023d) converter (part of the native ONNX codebase), when converting Keras to ONNX. FetaFix automatically identified the presence of a redundant Transpose node following the input node, and repaired it by (1) adjusting the Target model input dimensions so that they correspond to those of Source, (2) altering its tensor transposition order (permutation) property to align with the new input dimensions, and (3) updating the properties of nodes connected to Transpose to ensure proper tensor handling, reflecting the change in Transpose. This modification in particular was vital to case #$1414$ where FetaFix detected additional nodes (Gather, Unsqueeze, Add and Mul) preceding Transpose that apply input normalization. As part of the repair process, FetaFix updated the axes properties of Gather and Unsqueeze nodes to align with the updated input dimension and the Transpose permutation in Target.
@@
WB: Cases #$22$, $55$, $88$, and $1515$ were related to parameter precision discrepancies between Source and Target. Although the discrepancies were negligible, the model behavior deviated across the cases from 3.123.123.123.12-9.79.79.79.7%. FetaFix performed multiple iterations and managed to completely fix cases #$55$ and $1515$ while also significantly improving cases $22$ and $88$ by replacing the weights of Target with those of Source where differences were detected. We identified all cases to be related to the behavior and the setup of TFLiteConverter. The discrepancy after each repair step is depicted in Figure 4. At each repair step, FetaFix adjusts the weights & biases of a suspicious layer in Target with those from Source, in the order determined by the Layer Activation Analysis (Section 4.1.4), and evaluates the effects of the change by comparing the predictions of Source and Target.
@@
LH: From the $254254$ potential faults FetaFix localized, only $1212$ of them were actively affecting the model. These hyperparameters were associated with the pads hyperparameter of Conv and MaxPool nodes (cases #$1111$ & $1313$), as well as with the axes of Gather and Unsqueeze nodes responsible for input processing (case #$1414$). Repairing Target entailed adjusting their values with those from Source.
@@
CG: FetaFix identified $4747$ cases of graph-related potential faults. Of these, $44$ cases resulted in fault repairs that improved model performance (reduced output label discrepancy). The repairs played a crucial role in conversion cases #$1111$ and #$1313$ that initially presented discrepancies of 100100100100% and 35353535%, respectively. Our tool identified that, for case #$1111$, the converter had erroneously introduced multiple output nodes. In addition, three problematic padding nodes were found to affect the behavior of the models. FetaFix identified such faults and completely repaired the models by (1) removing any additional output nodes added to the model by the converter, and (2) repairing problematic padding nodes.
@@
For all the false positive cases observed in LH and CG, we performed manual inspection and observed that the converters performed semantically equivalent computation graph conversions. For instance, in the erroneous conversion presenting the largest number of CG fault localizations – ResNet101, Keras to TFLite (case #$88$) – Source contained Pad nodes which were absent in Target. However, Target contained a pads hyperparameter in the next convolutional layer that had the same effect. Also, BatchNorm nodes of Source were replaced by the converter with Mul and Add nodes in Target  effectively applying the normalization in a two-step semantically equivalent manner. FetaFix attempted to repair these scenarios by replacing them with the corresponding Source subgraph but these fixes had no effect on the number of output label discrepancies encountered.
@@
Figure 4. Model conversion cases that required an iterative weights and biases repair strategy with the percentage output label dissimilarity shown after each repair cycle (following preprocessing repair).
6.3. Fault Localization/Repair Effectiveness
@@
We consider effectiveness as the number of localized faults that could be repaired to improve Target prediction accuracy towards matching this of Source. We go through each fault category to determine the effectiveness of the fault localization component, as seen in the Total row in Table 2. We find that FetaFix had an 100100100100% efficacy rate (all potential fault occurrences detected could be repaired) for preprocessing (PP: 9/9999/99 / 9), input dimensions (ID: 4/4444/44 / 4) and tensor structure (TSS: 2/2222/22 / 2), as well as 98.198.198.198.1% efficacy for weights and biases (WB: 430/438430438430/438430 / 438). The weights and biases category had the most faults detected and repaired. FetaFix had much less success on cases related to the computation graph (CG), with $44$ of the $4747$ potential faults detected being repaired (8.58.58.58.5%) while other cases when repaired did not result in reduction of output label discrepancy. The CG fault repairs, although few in number, were instrumental in completely fixing conversion cases #$1111$ and $1313$. The effectiveness for LH category was low (4444%), with only $1212$ out of the $254254$ identified potential faults worth repairing. Following manual inspection, we infer that this happens because: (1) graph transformations across Source and Target may result in syntactic difference but be semantically equivalent. For instance, a BatchNorm node is transformed in Mul and Add nodes. Although this is localized as a potential fault by FetaFix  the repair employed by replacing Mul and Add nodes with BatchNorm will not reduce the output label discrepancy and is therefore rejected; (2) the LH difference detected is already mitigated by the converter by introducing additional nodes in the Target. For example, a pads (padding) hyperparameter in a Conv node on Source is replaced by a dominating Pad to the correspondent Conv node in target, leading to the same effect.
In terms of the fault analysis for layer-based types, we set FetaFix to perform fault localization with and without fault analysis, and we compared the outcomes. We concluded that FetaFix was able to repair the models in a similar manner with the results presented without the fault analysis, however enabling it applied fewer fault repairs, more effectively avoiding false positive cases where the applied repair was ineffective (up to 3.6% reduction). We also found the layer ranking assisted towards applying smaller repair steps, with the dissimilarity percentage decreasing more gradually with each repair cycle, making it easy to monitor and control the fault process.
6.4. Repair Validity
@@
In order to verify our results, we used the ImageNetV2 (Recht et al., 2019) dataset consisting of 10K images which was built with the purpose of providing test data for ImageNet, as an independent validation source. We checked the discrepancies of the repaired Target models against Source, and verified that (1) all cases that were fully repaired using ILSVRC2017 also presented no discrepancies with ImageNetV2, and (2) partially repaired cases #$22$ and $88$ presented 0.280.280.280.28% and 0.660.660.660.66% discrepancies, lower than those observed in ILSVRC2017. The full comparison can be found in Table 3. In addition, we used ImageNetV2 to verify that repaired models were giving consistent results against the dataset ground truths. Finally, we tested all $1010$ cases shown in Figure 1 presenting no discrepancies with ImageNetV2, to verify that no additional errors were detected. In detail, we verified in our experiments that the results of each repair cycle do not misclassify previously correct results, by testing all repaired models on ImageNetV2, checking the classification in two ways - against ground truth for the dataset inputs, and by comparing the results between source and target models. We did not observe any misclassification cases of correct Target model predictions prior to the repair process across the entirety of our experiment set.
6.5. Root Causes of Faults
We consider that many of the errors detected might have common root causes as part of the conversion process. Upon further manual examination, we observed that (1) all faults related with preprocessing could be repaired when the proper setting was selected, (2) errors related to Weights & Biases (WB) were related to an implicit action caused by tf2onnx converter, and (3) issues related to Input Dimensions (ID), Tensor Shape & Structure (TSS), Layer Hyperparameters (LH) and Computational Graph (CG) were attributed to problematic handling of the layers by the conversion tools (e.g., both torch.onnx and onnx2keras converters mishandled model inputs). We aim to report such issues to the development teams of each converter tool.
6.6. Results Generalizability
@@
We expect our approach to be applicable to any DL frameworks containing core computational layers and their conversion tools - including non-CNN models, requiring from the user to specify the core computational layers to be used in the comparison. In addition, the user must define a comparator for the evaluation of the model type under test (e.g., BLEU (Papineni et al., 2002) for text generation models). Providing proof of FetaFix effectiveness in this scenario, is subject to future work. In terms of our experiments, we chose image recognition models in our experiments as they are more tractable and need less resources than other DL architectures like transformers, as our experiments involved $1515$ model conversion cases with numerous iterations and analysis within each. All the models selected are widely used, either as-is or as the base for extension of models deployed on more complex tasks (e.g., ResNet101 extended for object detection). Our current implementation is bassed on ONNX, as it is a widely-used format, providing a comprehensive API for model definition and modification. However, our methodology can be easily extended to other representations.
@@
In terms of the fault types supported by FetaFix, they are based on a survey of issues encountered in discussion forums and empirical investigations (Table 1). There may be other fault types not currently considered by FetaFix. We are aware of this and we expect that FetaFix will keep evolving to handle emerging issues and other fault categories in model conversions between DL frameworks.
@@
Table 3. Experiments Overview: Discrepancies between Source and Target. before and after the repair process. for ILSVRC2017 and ImageNetV2 datasets.